OpenStudy (anonymous):

What is the exact value of the expressions the square root of 180. - the square root 125. + the square root of 5.? Simplify if possible. @punksauce

OpenStudy (anonymous):

@Catlover5925

OpenStudy (fibonaccichick666):

so, do you know how to factor 180?

OpenStudy (catlover5925):

4.472135955....I don't know what you round to

OpenStudy (anonymous):

What is the exact value of the expressions the square root of 180. - the square root 125. + the square root of 5.? Simplify if possible. 2the square root of 2. 12the square root of 2. 2the square root of 5. 12the square root of 5.

OpenStudy (anonymous):

its multiple choice

OpenStudy (fibonaccichick666):

you don't round, you need exact, also, @Catlover5925 please read http://openstudy.com/code-of-conduct

OpenStudy (fibonaccichick666):

anyways, @josh543 , can you factor 180?

OpenStudy (anonymous):

2*90

OpenStudy (anonymous):

You asked this exact same question and I responded to it. I guess you just ignored my help. That's cool.

OpenStudy (anonymous):

no

OpenStudy (jhannybean):

That's what he did to my response as well.

OpenStudy (anonymous):

im just not focused

OpenStudy (jhannybean):

Seems like this user asks for help and reposts to find an answer.

OpenStudy (fibonaccichick666):

ok, 2*90 is 180. so can you factor 90?

OpenStudy (anonymous):

*303

OpenStudy (anonymous):

3*30

OpenStudy (fibonaccichick666):

ok, so now factor 30

OpenStudy (anonymous):

3*10

OpenStudy (fibonaccichick666):

then factor 10.

OpenStudy (anonymous):

2*5

OpenStudy (fibonaccichick666):

cool, so now lets put that all together : 180=2*90 =2*3*30 =2*3*3*10 =2*3*3*2*5 Is this correct? Do you follow?

OpenStudy (anonymous):

wat shood i do with that

OpenStudy (fibonaccichick666):

1, when I ask if you understand please answer. 2, that is called a prime factorization, it is the easiest way to determine a square root.

OpenStudy (fibonaccichick666):

so again, do you understand?

OpenStudy (anonymous):

no

OpenStudy (fibonaccichick666):

which part?

OpenStudy (anonymous):

after doing the factors whats next

OpenStudy (fibonaccichick666):

if you don't understand the factoring we cannot continue, it is a vital concept.

OpenStudy (anonymous):

i already did the factors but what should i do with them

OpenStudy (fibonaccichick666):

but you just said you don't understand what we have done? If you understand, please find the prime factorization of 125.

OpenStudy (fibonaccichick666):

oh, and don't forget to show all of your work in finding that factorization

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5cube

OpenStudy (jhannybean):

Good job.

OpenStudy (fibonaccichick666):

how did you get that?